OpenStudy (anonymous):

Find cosθ if sinθ = -12/13 and tanθ > 0.

OpenStudy (anonymous):

A. -5/12 B. -5/13 C. 12/5 D. -13/12

OpenStudy (anonymous):

cos theta = 1-sin^2theta = 1 - (-12/13)^2

OpenStudy (anonymous):

costheta = \sqrt of the value u got

OpenStudy (anonymous):

\[\cos(x) = \sqrt{1-\sin^2(x)}\]

OpenStudy (anonymous):

How do I simplify that?

OpenStudy (anonymous):

sin(x)=-12/13 => \[\sin^2(x) = 12^2/13^2\]

OpenStudy (anonymous):

or you can use\[\sqrt{(1-\sin(x))(1+\sin(x))}\]

OpenStudy (anonymous):

But I don't understand how to get one of the answers.

OpenStudy (anonymous):

Where does the 5, 12, or 13 come in?

OpenStudy (anonymous):

[1+(12/13)] [1-(12/13)] multiply

OpenStudy (anonymous):

(25/13)(1/13)=(25/13^2)

OpenStudy (anonymous):

1 - (12/13) + (12/13) - (12/13)^2

OpenStudy (anonymous):

take the square root of that number

OpenStudy (anonymous):

you will get : \[\sqrt{\frac{ 5^2 }{13^2 }}\]

OpenStudy (anonymous):

So 5/13?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

But is it -5/13? Because there is no positive 5/13.

OpenStudy (anonymous):

yes its (-) because tan(x)>0

OpenStudy (anonymous):

Okay, thank you so much!! You are a lifesaver. :)

OpenStudy (anonymous):

\[\tan(x)=\frac{ \sin(x) }{ \cos(x) }\]

OpenStudy (anonymous):

if sin(x)<0 then MUST BE cos(x)<0

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

to make tan(x)>0

OpenStudy (anonymous):

Oh, I get it now. Thank you so much!

OpenStudy (anonymous):

any time